The IOs of 600 applicants to a certain college are approximately normally distributed with a mean of 115 and a standard deviation of 12. If the college requires an IQ of at least 95, how many of these students will be rejected on this basis, regardless of their qualifications? Note that IQs are recorded to the nearest integers.


1
Expert's answer
2022-03-30T16:28:46-0400

"P(X<95)=P(Z<\frac{95-115}{12})=P(Z<-1.67)=0.0475."

Number of the students rejected: "N=600*0.0475=29."